A 1.5 kg water bottle with potential energy of 17.9j. What is the height of the water bottle from ice surface
1
Expert's answer
2019-12-10T09:54:34-0500

Solution:


"E=m\times g \times H"

"m=1.5 kg"

"E=17.9 j"

"g=9.8 \frac {m}{s^2}"


"H=\frac {E}{mg}"

"H=\frac {17.9}{1.5\times 9.8}"


"H=1.22 m"

Answer: 1.22 m